### v1

@@ 4 Discussion
Our objective here has been to arrive at a small number of pairwise logratios of amalgamations (i.e., SLRs), where the amalgamations have substantive meaning for the practitioner (a biochemist, in this case) as well as being successful in explaining structural properties of the data set (the seasonal differences, in this case).
In a related work, Graeve and Greenacre (2020) performed PLR selection, also step-by-step, on the same data set, with a combination of statistical and domain-knowledge criteria (Greenacre, 2019). A list of PLRs that explained the highest percentages of logratio variance was determined statistically for the first step of the process. The biochemist selected one of these that had a clearly defined interpretation. This PLR was fixed and then a second list was prepared, that explained the next highest percentage of variance, for similar consideration by the biochemist. This process continued until six PLRs were chosen and deemed satisfactory. However, those six PLRs were probably quite specific to the data set. By contrast, in the present approach, the definition of the amalgamations can be considered more general and applicable as a benchmark set of variables. Using ratios of sums of parts is also more reliable than using ratios of single parts. Future studies of fatty acid compositions in marine biochemistry can use the set of seven amalgamation logratios (SLRs) proposed here (defined by the horizontal pairwise links between the amalgamations in Fig. 2) to see how well they represent the compositional data structure.
The overall conclusion is that the construction of a hierarchy of amalgamated parts, based on domain knowledge, and their transformation to logratios, is a useful alternative way of performing variable selection in compositional data analysis.

### v2

@@ 4 Discussion
Our objective has been to arrive at a small number of pairwise logratios of amalgamations (i.e., SLRs), where the amalgamations have substantive meaning for the practitioner (a biochemist, in this case) as well as being successful in explaining structural properties of the data set (the seasonal differences, in this case).
In a related work, Graeve:20 performed PLR selection, also step-by-step, on the same data set, with a combination of statistical and domain-knowledge criteria (Greenacre:19). A list of PLRs that explained the highest percentages of logratio variance was determined statistically for the first step of the process. The biochemist selected one of these that had a clearly defined interpretation. This PLR was fixed and then a second list was prepared, which explained the next highest percentage of variance, for similar consideration by the biochemist. This process continued until six PLRs were chosen and deemed satisfactory. However, those six PLRs were probably quite specific to the data set. By contrast, in the present approach, the definition of the amalgamations can be considered more general and applicable as a benchmark set of variables. Using ratios of sums of parts is also more reliable than using ratios of single parts. Future studies of fatty acid compositions in marine biochemistry can use the set of seven amalgamation logratios (SLRs) proposed here to see how well they represent the compositional data structure. The SLRs are defined by the horizontal pairwise links between the amalgamations in Fig. 2 and listed in Table 2 for future reference.
This idea of a hierarchy of amalgamations and their corresponding logratios can be exported to other fields, such as geochemistry where amalgamations of parts also occur naturally. For example, GreenacreGrunskyBaconShone:20 analyse 10 major oxides of the Aar massif in Switzerland, nine of which subdivide into three amalgamations: mafic, felsic and carbonate-apatite. Grunskyetal:24 study diamond-bearing geological formations and define three sets of amalgamations of 22-part cation compositions: crustal, mantle and kimberlite. There is also a classification of the parts into major, minor, rare and trace elements. There is a potential for subdivisions of these groups of elements into substantively meaningful subsets, as performed with the fatty acids in the present work.
The overall conclusion is that the construction of a hierarchy of amalgamated parts, based on domain knowledge of an expert, and their transformation to logratios, is a useful alternative way of performing variable selection in compositional data analysis. The success of this selection can be quantitatively assessed using concepts and measures particular to this statistical field.
Author declarations
The authors declare no competing interests. They also declare that they have used no AI tools in the preparation of this manuscript.
Data availability
The fatty acid data set in this paper can be retrieved from https://github.com/michaelgreenacre/CODAinPractice